OpenStudy (anonymous):

Please factor this: -n^2-3n+28/3n^3+9n^2-84n

OpenStudy (anonymous):

the first step is to factor out the n. What do you get?

OpenStudy (anonymous):

\[-n^2-3n+\frac{28}{3n^3}+9n^2-84n =0\] multiply this equation by 3n3 \[-3n^5-9n^3+28+27n^5-252n^4 =0\]

OpenStudy (anonymous):

Laura .. do complete or i do ?

OpenStudy (anonymous):

The whole thing is a fraction, by the way.

OpenStudy (anonymous):

\[24 n^5 -252 n^4 -9n^3+28 = zero\]

OpenStudy (anonymous):

Everything before the "/" is in the numerator, and everything after it is in the denominator.

OpenStudy (anonymous):

The part I don't know how to factor is 3n^3+9n^2-84n, which is in the denominator.

OpenStudy (anonymous):

Thank you for helping me with that.